OpenStudy (anonymous):

x + y = 5 y = one halfx + 2 If the two equations are graphed, at what point do the lines representing the two equations intersect? (2, 5) (5, 2) (2, 3) (3, 2)

OpenStudy (anonymous):

@angel12310 @CrazyTurtle483 @Destinyyyy

OpenStudy (anonymous):

I'm not good at these sorry. What grade r u in?

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

@ganeshie8 @Hayleymeyer @Keigh2015 @LynFran help

OpenStudy (hayleymeyer):

hmmm

OpenStudy (anonymous):

im also in 8th

OpenStudy (hayleymeyer):

wouldnt it be, B and C

OpenStudy (anonymous):

so a or d

OpenStudy (anonymous):

@Hayleymeyer

OpenStudy (hayleymeyer):

B and C

OpenStudy (anonymous):

so th answer is a or d?

OpenStudy (hayleymeyer):

u think its either A Or D?

OpenStudy (anonymous):

kinda

OpenStudy (hayleymeyer):

Go with D?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

thanks @Hayleymeyer

OpenStudy (lynfran):

rewrite the 1st equation we get y=5-x y=(1/2)x+2 5-x=(1/2)x+2 5-2=(1/2)x+x 3=(3/2)x 2=x when x=2 y=? y=5-x y=5-2 y=3 so C would be correct

OpenStudy (hayleymeyer):

@snozberry

OpenStudy (destinyyyy):

The correct answer is C.